# Optimised Storage for Datalog Reasoning

### 18 Dec 2023 v1

@@ Introduction
The computation and maintenance of materialisation have been well investigated. The standard seminaïve algorithm (Abiteboul, Hull, and Vianu 1995) efficiently computes the materialisation by avoiding repetitions during rule applications. This algorithm can also incrementally maintain the materialisation for fact additions. More general (incremental) maintenance algorithms like the counting algorithm (Gupta, Mumick, and Subrahmanian 1993), Delete/Rederive algorithm (Staudt and Jarke 1995), and Backward/Forward (B/F) algorithm (Motik et al. 2019) can maintain materialisation for both additions and deletions and are applicable beyond initial computation. Additionally, specialised algorithms optimised for particular rule patterns, like transitive closure (Subercaze et al. 2016), have been developed, and a modular framework proposed by Hu, Motik, and Horrocks (2022) combines standard approaches for normal rules with tailored approaches for specialised rules to further improve materialisation efficiency.
@@
Modular reasoning: Hu, Motik, and Horrocks (2022) present a modular version of the seminaïve algorithm, which integrates standard rule application with the optimised evaluation of specialised rules (e.g. transitive closure and chain rules). A Datalog evaluation is split into modules each of which manages a subset of the original program. The modular seminaïve algorithm is then obtained by replacing line 4 in Algorithm 1 with $A=A\cup(\Pi_{T}^{+}[I,\Delta]\setminus\Delta)$ for every module $T$, where $\Pi_{T}[I,\Delta]\setminus I\subseteq\Pi_{T}^{+}[I,\Delta]\subseteq\Pi_{T}^{\infty}[I\cup\Delta]\setminus I$.
@@
The above data structure can be computed by performing a post-order traversal starting from the root $a_{n}$. When a node is visited, its index is assigned by increasing the index of the previous node by one, and its interval is computed using the interval of its child. This requires only $O(n)$ time, as opposed to $O(n^{2})$ and $O(n^{3})$ required by existing approaches. In fact, for an arbitrary graph in general, the corresponding indexes and intervals can be constructed in $O(|V|+|E|)$ time, where $|V|$ and $|E|$ are the numbers of vertices and edges of the graph, respectively. In terms of space usage, in our particular example, the data structure requires $O(n)$ space, as opposed to $O(n^{2})$ required by a full materialisation. However, it should be noted that the worst-case space complexity is $O(|V|^{2})$. Space consumption can be reduced by choosing the optimum tree cover of a graph (Agrawal, Borgida, and Jagadish 1989), a technique that proves to be useful in our evaluation.

### 19 Dec 2023 v2

@@ Introduction
The computation and maintenance of materialisation have been well investigated. The standard seminaïve algorithm (Abiteboul, Hull, and Vianu 1995) efficiently computes the materialisation by avoiding repetitions during rule applications. This algorithm can also incrementally maintain the materialisation for fact additions. More general (incremental) maintenance algorithms like the counting algorithm (Gupta, Mumick, and Subrahmanian 1993), Delete/Rederive algorithm (Staudt and Jarke 1995), and Backward/Forward (B/F) algorithm (Motik et al. 2019) can maintain materialisation for both additions and deletions and are applicable beyond initial computation. Additionally, specialised algorithms optimised for particular rule patterns, like transitive closure (Subercaze et al. 2016), have been developed, and a modular framework proposed by Hu, Motik, and Horrocks (2022) combines standard approaches for normal rules with tailored approaches for certain types of rules to further improve materialisation efficiency.
@@
Modular reasoning: Hu, Motik, and Horrocks (2022) present a modular version of the seminaïve algorithm, which integrates standard rule application with the optimised evaluation of certain rules (e.g. transitive closure and chain rules). A Datalog evaluation is split into modules each of which manages a subset of the original program. The modular seminaïve algorithm is then obtained by replacing line 4 in Algorithm 1 with $A=A\cup(\Pi_{T}^{+}[I,\Delta]\setminus\Delta)$ for every module $T$, where $\Pi_{T}[I,\Delta]\setminus I\subseteq\Pi_{T}^{+}[I,\Delta]\subseteq\Pi_{T}^{\infty}[I\cup\Delta]\setminus I$.
@@
The above data structure can be computed by performing a post-order traversal starting from the root $a_{n}$. When a node is visited, its index is assigned by increasing the index of the previous node by one, and its interval is computed using the interval of its child. This requires only $O(n)$ time, as opposed to $O(n^{2})$ and $O(n^{3})$ required by existing approaches. In terms of space usage, in our particular example, the data structure requires $O(n)$ space, as opposed to $O(n^{2})$ required by a full materialisation. For an arbitrary graph in general, the corresponding indexes and intervals can be constructed in $O(|V|+|E|)$ time, where $|V|$ and $|E|$ are the numbers of vertices and edges of the graph, respectively, and the worst-case space complexity is $O(|V|^{2})$. Note that space consumption can sometimes be reduced by choosing the optimum tree cover of a graph (Agrawal, Borgida, and Jagadish 1989), a technique that proves to be useful in our evaluation.